# Flow-Anchored Consistency Models

### v1

@@ 3.2.3 FACM Training Algorithm
With the objective functions and conditioning signals defined, we now present the complete FACM training strategy in Algorithm 1, which illustrates how the FM and CM losses are computed and combined in a single training step. The computation of the total derivative $\nabla_{t}\bm{F}_{\theta}$ in the CM loss is performed using a Jacobian-vector product, denoted as JVP. The “tangents” are given by the tuple $(\bm{v},1)$, where $\bm{v}$ is the tangent with respect to the input $\bm{x}_{t}$, and $1$ is the tangent for the conditioning $c_{\text{CM}}$. We provide a formal proof in Appendix A.1 for why the condition tangent is always 1.
In summary, the principle of Flow-Anchoring offers a robust and fundamental solution. While other methods achieve stability, they do so implicitly and less effectively. For instance, the standard CM objective, when stabilized by methods like sCM (Lu & Song, 2024), can eventually converge to describe the velocity field, but this is an indirect consequence rather than a direct objective. Other approaches like MeanFlow (Geng et al., 2025) implicitly retain this capability by modeling the average velocity to arbitrary endpoints. However, this generalization creates an overly-coupled learning problem. Because learning the instantaneous velocity is just one special case within this broader objective, the two distinct tasks become entangled in a single loss, leading to optimization difficulties and sub-optimal performance. In contrast, FACM is a more direct and principled solution to ensure the model maintains a stable representation of the flow throughout training.
4 Experiments
4.1 Experimental Setup
We empirically validate FACM on image generation benchmarks, including CIFAR-10 (Krizhevsky & Hinton, 2009) and ImageNet 256$\times$256 (Deng et al., 2009). We evaluate models based on Fréchet Inception Distance (FID) (Heusel et al., 2017) and the Number of Function Evaluations (NFE). FACM can be trained from scratch or by distilling a pre-trained model. Our default experimental setup involves a two-stage process. We first pre-train a standard FM model, incorporating our mixed-objective conditioning as detailed in Appendix A.3 (a) to accelerate the subsequent distillation. We then distill this teacher using the FACM strategy. For few-step inference, we follow the standard multi-step sampling procedure for CMs as described in Song et al. (2023) and Appendix A.3 (b). Further details on our experimental settings are provided in Appendix A.3.
4.2 Main Results
4.2.1 Comparison with State-of-the-art
As shown in Table 1, FACM achieves state-of-the-art results on both CIFAR-10 and ImageNet 256$\times$256. Specifically, our method achieves FIDs of 1.76 (NFE=1) and 1.32 (NFE=2) on ImageNet 256$\times$256 by training a LightningDiT model in latent-space, and 2.69 (NFE=1) and 1.87 (NFE=2) on CIFAR-10 by training a DDPM++ model (Ho et al., 2020) in pixel-space, significantly outperforming previous methods on both benchmarks. Remarkably, our few-step model even surpasses some multi-step baselines that require hundreds of function evaluations.
@@ 4.2.1 Comparison with State-of-the-art
Table 1: Few-step generation on CIFAR-10 and ImageNet 256$\times$256. “$\times$2” indicates that CFG doubles the NFE per step. Our method sets a new state-of-the-art on both datasets.
@@
iCT (Song & Dhariwal, 2023) 1 2.83
eCT (Geng et al., 2024) 1 3.60
sCM (sCT) (Lu & Song, 2024) 1 2.85
IMM (Zhou et al., 2025) 1 3.20
MeanFlow (Geng et al., 2025) 1 2.92
@@
TRACT (Berthelot et al., 2023) 2 3.32
CD (LPIPS) (Song et al., 2023) 2 2.93
iCT-deep (Song & Dhariwal, 2023) 2 2.24
ECT (Geng et al., 2024) 2 2.11
sCM (sCT) (Lu & Song, 2024) 2 2.06
IMM (Zhou et al., 2025) 2 1.98
@@
Class-Conditional ImageNet 256$\times$256
@@
SiT-XL/2 (Ma et al., 2024) 675M 250$\times$2 2.06
DiT-XL/2 (Peebles & Xie, 2023) 675M 250$\times$2 2.27
REPA (Yu et al., 2025) 675M 250$\times$2 1.42
LightningDiT (Yao et al., 2025) 675M 250$\times$2 1.35
@@
iCT (Song & Dhariwal, 2023) 675M 1 34.24
Shortcut (Frans et al., 2025) 675M 1 10.60
MeanFlow (Geng et al., 2025) 676M 1 3.43
FACM (Ours) 675M 1 1.76
@@
iCT (Song & Dhariwal, 2023) 675M 2 20.30
IMM (Zhou et al., 2025) 675M 1$\times$2 7.77
MeanFlow (Geng et al., 2025) 676M 2 2.20
@@ 4.3 Ablation Study on the Training Strategy
We conduct ablation studies to validate our claims regarding the training strategy. We test on the ImageNet 256$\times$256 dataset by distilling a pre-trained LightningDiT model. The results provide strong evidence for our central claim: the presence of the FM objective is the critical stabilizing anchor.
Table 2: FID scores (NFE=2) on ImageNet 256$\times$256 for different few-step methods applied to various multi-step FM models. $\dagger$ indicates our reproduction.
Method Baseline (NFE=250$\times$2) sCM MeanFlow FACM (Ours)
@@
Different Architectures. To demonstrate the architectural agnosticism of our approach, we apply FACM, sCM, and MeanFlow to a range of state-of-the-art architectures, including SiT-XL/2, REPA, DiT-XL/2, and LightningDiT. All methods are distilled from their respective multi-step FM models. As shown in Table 2, FACM consistently achieves the lowest FID scores across all tested backbones. This highlights that Flow-Anchoring is a fundamental principle for stabilizing consistency training that is not limited to a specific model design. Notably, methods like REPA and LightningDiT, which improve performance by applying distribution constraints to the VAE, retain these advantages in the few-step generation setting.
Stabilization Strategy. To ensure a fair comparison between different stabilization methods for few-step generative models, we conduct a controlled experiment comparing sCM, MeanFlow, and our FACM. This setup isolates the effectiveness of each strategy by distilling from an identical, pre-trained LightningDiT teacher model where possible. As official code was unavailable, we reproduced the compared methods following their original descriptions (Detailed in Appendix A.3 (c).). Architectural modifications integral to these methods were handled as follows: for sCM, we pre-trained the student model with the required pixel normalization layer from the start; for MeanFlow, the necessary additional time embedder was zero-initialized and incorporated during the distillation phase. Our FACM requires no architectural changes.
As shown in Table 3, while all methods achieve stable training, our FACM framework provides a more fundamental solution and superior results. The sCM distillation without pixel norm (cannot be added post-hoc via fine-tuning) collapses as predicted. While sCM incorporates various stabilization techniques, it only achieves a limited and fragile training stability; we found that reducing the tangent warmup steps or altering the time sampling schedule can lead to randomly occurring training collapse. MeanFlow achieves robust stability through its unified objective, but we diagnose that its conceptual unification of instantaneous and average velocity ($u(z,t,t)=v(z,t)$) leads to an over-coupling of the two tasks, which hinders optimization and slows convergence. In contrast, FACM explicitly separates the two tasks, allowing it to more effectively leverage the velocity field provided by the pre-trained FM teacher as a flow anchor while stably learning the shortcut objective.
Weighting Functions. We analyze the CM loss weighting functions (Table 4), which are crucial for navigating the trade-off between ensuring endpoint quality (high-SNR regions) and satisfying global consistency (low-SNR regions). We find that $1-t^{p}$ and $\cos(t\cdot\pi/2)$ serve as effective general weighting solutions. In our experiments, $p=0.5$ yielded the best results, appropriately focusing learning on the more informative low-SNR regime.
Impact of Teacher Model Convergence. To further validate our Flow-Anchoring hypothesis, we compare how sCM and our FACM perform when distilled from teacher FM models at different stages of pre-training. Figure 3(b) shows a striking divergence. The performance of sCM degrades when using a highly-converged teacher, which implies a fragile training process that requires an empirical trade-off: one must find a moderately-converged teacher, as a better flow model does not guarantee a better shortcut model. This is because the more complex flow of a highly-converged teacher is harder to model implicitly through the shortcut objective alone. In contrast, FACM’s performance scales monotonically with the teacher’s quality. This is a critical finding: Flow-Anchoring removes this empirical guesswork, transforming the two-stage training from a delicate balancing act into a principled optimization where a better FM model directly yields a better CM model.
Table 3: Ablation on stabilization strategies. All methods are distilled from the same LightningDiT teacher (architectural modifications are only made to the student model). $\dagger$: Our reproduction.
Method Params Pre-train epochs Distill epochs FID (NFE=1, $\downarrow$) Stable
@@
sCM (w/ pixel norm.) 676M 800 250 3.04 $\checkmark$
MeanFlow 676M 800 250 2.75 $\checkmark$
FACM (Auxiliary Condition) 676M 800 250 1.93 $\checkmark$
FACM (Expanded Interval) 675M 800 250 1.76 $\checkmark$
MeanFlow (from scratch) 676M 0 1120 2.65 $\checkmark$
FACM (from scratch) 675M 0 800 2.27 $\checkmark$
4.3.1 Discussion: From-Scratch Training vs. Distillation
While our method can be trained from scratch and achieves a competitive result (See Table 3), we identify the two-stage distillation paradigm as the more principled and practically superior approach. Attempting to learn both the anchor and the shortcut simultaneously from scratch introduces a “chicken-and-egg” problem, as the model must learn a shortcut based on a trajectory it has not yet accurately modeled. This creates an unstable “moving target” for optimization and incurs higher computational costs. In contrast, distillation from a pre-trained FM teacher provides a fixed, high-quality velocity field, offering a much more stable and well-defined learning objective. MeanFlow (Geng et al., 2025) also encounters this problem in the from-scratch setting, achieving its optimal performance only by having its objective degenerate to a Flow Matching task for a large portion of samples (e.g., 75%), which further validates our core thesis that a robust foundation in the velocity field is a prerequisite for learning stable shortcuts.
Table 4: Ablation on weighting functions for the CM loss component on ImageNet (NFE=2).
Weighting $\alpha(t)$ $\beta(t)$ FID ($\downarrow$) Weighting $\alpha(t)$ $\beta(t)$ FID ($\downarrow$)
$1.0$ $1.0$ 1.45 $1-t^{0.5}$ $1-t^{0.5}$ 1.37
$1.0$ $\cos(t\cdot\pi/2)$ 1.39 $\cos(t\cdot\pi/2)$ $\cos(t\cdot\pi/2)$ 1.40
$1-t^{0.5}$ $1.0$ 1.42 $1-t^{0.5}$ $\cos(t\cdot\pi/2)$ 1.32
Figure 3: (a) Weighting functions used in the CM loss. (b) Performance of distilled models (NFE=2) vs. FM pre-training epochs. (c) Performance of distilled models vs. FACM distillation epochs.

### v2

@@ 3.3.3 Training Algorithm and Scalable Chain-JVP Implementation
With the objective functions and conditioning signals defined, we present the complete FACM training strategy in Algorithm 1. A key component of this algorithm is the computation of the total derivative $\nabla_{t}\bm{F}_{\theta}$ in the CM loss (Line 7), performed using a Jacobian-vector product (JVP).
The JVP computation, however, presents critical bottlenecks when using modern acceleration techniques. While its incompatibility with components like Flash Attention 2 (dao2023flashattention2) can be resolved using methods from sCM (lu2024simplifying), a more fundamental memory bottleneck emerges from its conflict with Fully Sharded Data Parallel (FSDP) (zhao2023pytorchfsdpexperiencesscaling). Standard JVP implementations require the model’s full parameters $\theta$ to be materialized on the device, forcing an all_gather operation in an FSDP setup. This reconstructs the entire parameter set on each GPU, causing a prohibitive memory spike that makes training models with over ten billion (10B) parameters impossible. To overcome this, we leverage the chain rule. For a network composed of modules $\bm{F}_{\theta}=f_{L}\circ\dots\circ f_{1}$, the JVP can be computed sequentially:
$J_{\bm{F}_{\theta}}(\bm{z})\cdot\bm{v}=J_{f_{L}}(\bm{z}_{L-1})\cdot(\dots\cdot(J_{f_{2}}(\bm{z}_{1})\cdot(J_{f_{1}}(\bm{z}_{0})\cdot\bm{v}))\dots)$ (15)
where $\bm{z}_{i}=f_{i}(\bm{z}_{i-1})$ is the intermediate output. Our approach computes the JVP for each module sequentially, embedding this operation within the FSDP logic. Its speed is consistent with a standalone JVP pass, adding only standard FSDP overhead. This ensures that only one module’s parameters are materialized at a time. Consequently, peak memory depends on the largest module, not the entire model, and the resulting memory savings grow with the model’s parameter count.
In summary, the principle of Flow-Anchoring offers a robust and fundamental solution. While other methods achieve stability, they do so with certain limitations. For instance, sCM (lu2024simplifying) requires architectural modifications to its normalization layers, limiting its adaptability to large, pre-trained models. Other approaches like MeanFlow (geng2025meanflowsonestepgenerative), while clever, present a trade-off: by treating the instantaneous velocity as merely an edge case ($r=t$) of the primary average velocity objective, the learning tasks become over-coupled. As a result, the supervisory signal for the underlying flow is often diluted, which we have observed can lead to training collapses and underfitting. In contrast, FACM provides a more direct and principled solution. Through our innovative expanded time interval strategy, the anchoring and shortcut tasks are functionally decoupled into distinct domains. This ensures the flow anchor receives a clear, undiluted supervisory signal at all times, forcing the model to maintain a stable and high-fidelity representation of the flow. This robust theoretical foundation, combined with our scalable Chain-JVP implementation, makes FACM not only stable but also highly practical for training models at an unprecedented scale.
@@
Table 1: Few-step generation on CIFAR-10 and ImageNet $\times256$. $\times$2” indicates that CFG doubles the NFE per step. Our method sets a new state-of-the-art on both datasets.
@@
iCT (song2023improved) 1 2.83
eCT (geng2024consistency) 1 3.60
sCM (sCT) (lu2024simplifying) 1 2.85
IMM (zhou2025inductive) 1 3.20
MeanFlow (geng2025meanflowsonestepgenerative) 1 2.92
@@
TRACT (berthelot2023tract) 2 3.32
CD (LPIPS) (song2023consistency) 2 2.93
iCT-deep (song2023improved) 2 2.24
ECT (geng2024consistency) 2 2.11
sCM (sCT) (lu2024simplifying) 2 2.06
IMM (zhou2025inductive) 2 1.98
@@
Class-Conditional ImageNet $\times256$
@@
[…]
@@
iCT (song2023improved) 675M 1 34.24
Shortcut (frans2025one) 675M 1 10.60
MeanFlow (geng2025meanflowsonestepgenerative) 676M 1 3.43
FACM (Ours) 675M 1 1.70
@@
iCT (song2023improved) 675M 2 20.30
[…]
MeanFlow (geng2025meanflowsonestepgenerative) 676M 2 2.20
@@
4 Experiments
4.1 Experimental Setup
We empirically validate FACM on image generation benchmarks, including CIFAR-10 (Krizhevsky09) and ImageNet $\times256$ (deng2009imagenet). We evaluate models based on Fréchet Inception Distance (FID) (heusel2017gans) and the Number of Function Evaluations (NFE). FACM can be trained from scratch or by distilling a pre-trained model. Our default experimental setup involves a two-stage process. We first pre-train a FM model, incorporating our mixed-objective conditioning as detailed in Appendix A.4 (a) to accelerate the subsequent distillation. We then distill this teacher using the FACM strategy. For few-step inference, we follow the standard multi-step sampling procedure for CMs as described in Appendix A.4 (b). Further details on our experimental settings are provided in Appendix A.4. To demonstrate scalability, we also distill a 14B parameter model (Wan2.2) on the text-to-image (T2I) task, achieving high-fidelity generation in just 2-8 steps.
4.2 Main Results
4.2.1 Comparison with State-of-the-art
As shown in Table 1, FACM achieves state-of-the-art results on both CIFAR-10 and ImageNet $\times256$. Specifically, our method achieves FIDs of 1.70 (NFE=1) and 1.32 (NFE=2) on ImageNet $\times256$ by training a LightningDiT model in latent-space, and 2.69 (NFE=1) and 1.87 (NFE=2) on CIFAR-10 by training a DDPM++ model (ho2020denoising) in pixel-space, significantly outperforming previous methods on both benchmarks. Remarkably, our few-step model even surpasses some multi-step baselines that require hundreds of function evaluations.
@@ 4.3 Ablation Study on the Training Strategy
We conduct ablation studies to validate our claims regarding the training strategy. We test on the ImageNet $\times256$ dataset by distilling a pre-trained LightningDiT model. The results provide strong evidence for our central claim: the presence of the FM objective is the critical stabilizing anchor.
Table 2: FID scores (NFE=2) on ImageNet $\times256$ for different few-step methods applied to various backbone architectures. $\dagger$ indicates our reproduction.
[…]
@@
Table 3: Ablation on stabilization strategies. All methods are distilled from the same LightningDiT teacher. $\dagger$: Our reproduction. $\ast$: For sCM, more epochs yield worse results.
Method Params FM epochs CM epochs FID (NFE=1, $\downarrow$) Stable
@@
sCM (w/ pixel norm.) 676M 600 30 3.04 $\checkmark$
MeanFlow 676M 800 200 2.75 $\checkmark$
FACM (Auxiliary Condition) 676M 800 200 1.97 $\checkmark$
FACM (Expanded Interval) 675M 800 200 1.81 $\checkmark$
Training from scratch methods
MeanFlow 676M 0 1120 2.65 $\checkmark$
FACM (Expanded Interval) 675M 0 800 2.27 $\checkmark$
Different Architectures. To demonstrate the architectural agnosticism of our approach, we apply FACM, sCM, and MeanFlow to a range of state-of-the-art architectures, including SiT-XL/2, REPA, DiT-XL/2, and LightningDiT. All methods are distilled from their respective multi-step FM models. As shown in Table 2, FACM consistently achieves the lowest FID scores across all tested backbones. This highlights that Flow-Anchoring is a fundamental principle for stabilizing consistency training that is not limited to a specific model design.
Stabilization Strategy. To ensure a fair comparison, we distill sCM, MeanFlow, and FACM from an identical LightningDiT teacher (reproduction details in Appendix A.4 (c)). As shown in Table 3, FACM achieves superior results due to its principled approach to stability without requiring architectural changes. In contrast, sCM’s stability is limited, depending on architectural modifications (pixel normalization) and sensitive hyperparameter tuning. MeanFlow achieves robustness but at the cost of an over-coupled objective ($u(z,t,t)=v(z,t)$) that hinders optimization by diluting the essential path modeling task. FACM’s explicit task separation proves more effective, as it allows the model to stably learn the shortcut while remaining anchored to the teacher’s flow.
Sensitivity to Teacher Model Quality. As shown in Figure 4(a), FACM’s performance monotonically improves with teacher quality. This demonstrates that by explicitly anchoring the teacher’s complex flow, our method can consistently benefit from stronger teachers. This confirms FACM acts as a high-fidelity trajectory compression rather than a lossy compromise on the pre-trained flow.
Ablation on Key Components. As shown in Table 4, introducing Flow-Anchoring with our Expanded Time Interval decouples the FM and CM tasks, yielding faster convergence. Fidelity is further improved as shortcut interpolation ($\alpha$) and beta weighting ($\beta$) ensure a smooth transition to FM supervision as $t\to 1$, while residual clamping suppresses gradient spikes. Together, these components stably guide the learning dynamic via the FM anchor, leading to significantly better trajectory fidelity.
Table 4: Ablation study on key techniques on ImageNet 256x256.
Configuration FID@Epochs 10 (NFE=1, $\downarrow$) Collapse
MeanFlow / Fixed $r=1$ MeanFlow (0% FM) 372.3-391.5 Yes
MeanFlow (75% FM) 43.03 No
Fixed $r=1$ MeanFlow (75% FM) 15.54 No
w/ Flow Anchoring (Expanded Time Interval) 4.31 No
w/ Interpolation ($\alpha(t)=1$) 3.42 No
w/ Residual Clamping 2.86 No
w/ Beta Weighting ($\beta(t)=1$) (FACM) 2.51 No
Sensitivity to FM Loss Weight. The FM loss is a prerequisite for stability, but the minimum required weight $\lambda_{\text{FM}}$ depends on the model’s initialization. Our investigation reveals a nuanced picture that strongly supports a simple default choice (e.g., $\lambda_{\text{FM}}=1.0$). As summarized in Table 5 (left vs. right), the non-finetuned setting requires at least $\lambda_{\text{FM}}\geq 0.1$ to avoid collapse, whereas the finetuned setting remains stable with $\lambda_{\text{FM}}$ as low as $10^{-8}$. These results lead to a key conclusion: while a non-zero FM weight is essential, FACM is highly robust to the specific weight across several orders of magnitude once stability is achieved. This robustness, which stems from our decoupled design, makes a direct summation a simple, effective, and reliable choice that avoids costly hyperparameter tuning.
Table 5: Sensitivity to $\lambda_{\text{FM}}$ under two settings. Left: model not pre-finetuned on $1<t<2$. Right: model is pre-finetuned on $1<t<2$.
FM Loss Weight ($\lambda_{\text{FM}}$) FID (NFE=1, $\downarrow$)
0.0–0.1 Collapse
0.1–10.0 3.17–3.22
10.0–64.0 3.32–4.97
FM Loss Weight ($\lambda_{\text{FM}}$) FID (NFE=1, $\downarrow$)
0.0-1e-8 Collapse
1e-8–1e-4 2.90-5.88
1e-4–10.0 2.90–3.02
10.0–64.0 3.02–4.58
4.3.1 Training Dynamics of FACM
We analyze the training dynamics by plotting the total gradient norm under different configurations in Figure 3. Figure 3(a) clearly shows that removing the FM objective leads to catastrophic gradient spikes, after which the model’s output immediately degenerates into pure noise (mode collapse). This confirms our hypothesis that a pure consistency gradient can trap the model in a local optimum where it sacrifices endpoint fidelity in pursuit of global consistency. Figure 3(b) further illustrates the effect of our auxiliary techniques. While each individually helps to suppress the gradient norm compared to removing them all, their combined use in our baseline model achieves the lowest and most stable gradient profile, demonstrating their synergistic effect in stabilizing the training process.
Figure 3: (a) The raw gradient norm for a pure CM (w/o FM Objective) shows an instantaneous spike leading to collapse, while our baseline remains stable. (b) The smoothed gradient norm for ablations of auxiliary techniques. Removing any single technique increases instability.
4.3.2 Scalability on a 14B Text-to-Image Model
Scaling continuous-time consistency models to billion-parameter scales presents a significant challenge due to the Jacobian-vector product (JVP) computation. While recent Differential Derivation Equation (DDE) sun2025unified; wang2025transition, can yield results comparable to JVP on models up to 1B parameters, we observe that they exhibit significant deviation on larger models like our 14B setup. In such cases, their computed derivatives become nearly orthogonal or even opposed to the JVP result, indicating an inherent error accumulation that hinders further scalability. To address this, our memory-efficient Chain-JVP provides an accurate and scalable solution. To demonstrate its effectiveness, we applied FACM to distill the 14B parameter Wan 2.2 model. This process successfully accelerated inference from $2\times 40$ steps to just 2-8 steps. For the experiment, we used a pre-trained Wan 2.2 Text-to-Video (T2V) model (wan2025) as a teacher on an in-house Text-to-Image (T2I) dataset (Despite being a T2V model, Wan 2.2 has strong image generation capability from mixed image-video pre-training.). Furthermore, we adapt the model’s self-attention and cross-attention mechanisms to be compatible with JVP computation, following the formulation of (lu2024simplifying). This adaptation also addresses the correctness of differentiation for variable-length sequences and with bf16 precision. Our visualizations for this experiment are provided in Figure 1 and Appendix A.10, including comparison against the baseline model, as well as the FLUX.1-Dev (flux2024_dev) and the FLUX.1-Schnell models (flux2024_sch).
4.4 From Consistency Models to Flow Mapping Models
Recent work has increasingly emphasized the advantages of Flow Mapping, where the model learns to predict the average velocity from an arbitrary time $t$ to another time rr(sabour2025align; wang2025transition; geng2025meanflowsonestepgenerative; boffi2025flowmapmatchingstochastic). Flow Mapping requires the model to ensure that the derivative of $f_{\theta}(x_{t},t,r)=x_{t}+(r-t)F_{\theta}(x_{t},t,r)$ is zero. This formulation is an extension of consistency models along the trajectory, demanding that the model’s prediction, $f_{\theta}(x_{t},t,r)$, remains consistent over any time interval $[0,r]$ (detailed in Appendix A.6). We found through experiments on Wan2.2 that FACM can be easily adapted to be compatible with the Flow Mapping formulation. This is achieved simply by changing $c_{\text{CM}}$ from $(t)$ to $(t,r)$ through zero-initialized time embedder and projection modules, while $c_{\text{FM}}$ is maintained in the separate, expanded time domain. As illustrated in Figure 4(c), the Expanded Time Interval strategy allows the Flow Mapping to be more stably anchored to the teacher’s trajectory. If the prediction of the instantaneous velocity field is treated merely as a marginal case of the Auxiliary Time Condition (e.g., $r=t$), the FM loss becomes highly unstable, even when increasing the sampling proportion of $t=r$ as is done in MeanFlow. The consistently lower FM loss for the FACM condition demonstrates the superior trajectory fidelity achieved by our decoupled training strategy.
Figure 4: (a) Performance of student models (NFE=2) vs. teacher FM model pre-training epochs. (b) Performance vs. distillation epochs. (c) Trajectory fidelity analysis of 14B Wan2.2 model via flow matching loss. Apart from the conditioning method, all other settings were the same.